# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

### v1

@@ 1. We prove the theorem by establishing that, under the assumption:
$({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}+{\mathbf{v}}_{\varepsilon}+\nabla(i(\rho))_{\varepsilon}+{\bf r}^{(2)}_{\varepsilon}=0,$ (3.26)
@@
$r^{(1)}_{\varepsilon}=(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}),\qquad{\bf r}^{(2)}_{\varepsilon}=\big{(}({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big{)}_{\varepsilon}-({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}.$
@@
${\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}+(1+\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon})\omega_{\varepsilon}+\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=0\qquad\,\,\mbox{in $\Omega^{r}$}.$ (3.27)
@@
$=\rho_{\varepsilon}X^{(\varepsilon)}+r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}.$
@@
${\mathcal{R}}_{\varepsilon}=f^{\prime}(X^{(\varepsilon)})\big{(}r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\big{)}-f(X^{(\varepsilon)})r^{(1)}_{\varepsilon}.$ (3.30)
@@
Let $\zeta\in C^{\infty}(\mathbb{R}^{2})$ satisfy (3.21). Multiply the last equation by $\zeta$, integrate over $\Omega$, and integrate by parts via using (3.21) and (3.24) to obtain
@@
${\mathcal{R}}_{\varepsilon}\to 0\qquad\mbox{in $L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.34)
@@
$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\to 0\qquad\,\mbox{in$L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.36)
@@
In the argument below, we use (3.23). We see that $\nabla(\rho_{\varepsilon}-\rho)\cdot{\mathbf{v}}_{\varepsilon}\to 0$ in $L^{2}(\Omega^{r})$ as $\varepsilon\to 0$, because $\nabla\rho_{\varepsilon}\to\nabla\rho$ in $L^{2}(\Omega^{r})$ and ${\mathbf{v}}_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$. Also, $(\rho_{\varepsilon}-\rho)(\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}-\mbox{\rm div}\,{\mathbf{v}})\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and $\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}\to\mbox{\rm div}\,{\mathbf{v}}$ in $L^{2}(\Omega^{r})$. Finally, $(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and converge to zero a.e. in $\Omega_{r}$, and $\mbox{\rm div}\,{\mathbf{v}}\in L^{2}(\Omega^{r})$, so that
@@
$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=\partial_{\xi_{1}}\big{(}({\mathbf{v}}\cdot\nabla v_{2})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{2})_{\varepsilon}\big{)}-\partial_{\xi_{2}}\big{(}({\mathbf{v}}\cdot\nabla v_{1})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{1})_{\varepsilon}\big{)}.$
Then (3.36) follows from Lemma A.1 in Appendix A with $p=q=2$, $b=v_{j}$, and $u=\partial_{j}v_{k}$ for the corresponding $j,k=1,2$, and $i=3-k$.
@@
where ${\mathbf{v}}^{-}=(v_{1}^{-},v_{2}^{-})$ and ${\mathbf{v}}=(v_{1},v_{2})$. Taking the tangential derivative $\partial_{\boldsymbol{\tau}}:=\partial_{\xi_{1}}+f_{s}^{\prime}\partial_{\xi_{2}}$ of the Rankine-Hugoniot conditions (3.42) along $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and using the condition that $f_{s}^{\prime}=0$ at $P$, we obtain that, at point $P$,
@@
$00=-\int_{\Gamma_{\rm shock}}\big{|}\frac{\omega}{\rho}\big{|}^{2}\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}\,{\rm d}l>$ 0 .
@@
This completes the proof of Theorem 3.2. ∎
@@ 4. Applications to Transonic Shock problems
Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem which satisfies conditions (i)–(vii) of Definition 3.1 and assumption (i) of Theorem 3.2. Assume that
@@
    $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is not a segment of straight line. In particular, denoting by $P_{1}$ and $P_{2}$ the endpoints of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, there exists a point $P^{*}\in\overline{\Gamma_{1}^{int}}\setminus\{P_{1}\}$ such that $\boldsymbol{\tau}(P_{1})\neq\pm\boldsymbol{\tau}(P^{*})$, where $\boldsymbol{\tau}(\cdot)$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at a point.
@@
Therefore, the remaining possibility is that $\hat{Q}=Q$. Note that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(Q)=0$. Moreover, using condition (c), the Rankine-Hugoniot conditions, and the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and of $(\rho,{\mathbf{v}})$ given in condition (v) of Definition 3.1 and assumption (i) of Theorem 3.2, we see that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P_{1})=({\mathbf{v}}\cdot\boldsymbol{\tau}_{S})(P_{1})\neq 0$. This implies that $Q\neq P_{1}$. Also, since $\hat{Q}\neq P^{*}$ as we discussed above, then $Q\neq P^{*}$ for the present case. Thus, for the present case $Q=\hat{Q}$, it follows that $Q$ is an interior point of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, and the part of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ between $P_{1}$ and $Q$ lies on the straight line $S$. In particular, line $S$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q$. We now shift and rotate the coordinates to have the origin at $Q$ and the coordinates $\xi_{1}$ and $\xi_{2}$ to be along $S$ and $L$, respectively. Then $O_{1}=(0,\tilde{v}_{1})$ for some $\tilde{v}_{1}\in\mathbb{R}$ (in fact, $\tilde{v}_{1}\neq 0$ by condition (v) of Definition 3.1, but this will not be used below). To fix notation, let the $\xi_{1}$–axis along $S$ be oriented so that $P_{1}=(\xi_{P_{1}},0)$ with $\xi_{P_{1}}<0$. Since $S=\{\xi_{2}=0\}$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q=(0,0)$, then curve $\Gamma^{\mbox{\rm\footnotesize int}}_{1}\cap B_{r}(Q)$ is a graph for some $r>0$: There exists $f\in C^{2}(\mathbb{R})$ such that
@@
$({\mathbf{v}}_{1}\cdot\boldsymbol{\tau})(P)=\frac{-\xi_{1}+(\tilde{v}_{2}-f(\xi_{1}))f^{\prime}(\xi_{1})}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}=\frac{-\xi_{1}+O(\varepsilon)\xi_{1}}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}\neq 0\qquad\mbox{for $\xi_{1}\in(0,\varepsilon)$ if $\varepsilon$ is small.}$.
@@ 4.1. Lower Regularity of the Regular Shock Reflection Solutions for the Isentropic Euler System
[…]
where +$\mathbb{R}^{2}_{+}=\mathbb{R}^{2}\cap\{x_{1}>0\}$. We seek functions $(\rho,{\mathbf{u}})(\mathbf{x},t)$ satisfying system (2.1) in $\Lambda$ with the boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on $\partial\Lambda$ and the initial data:
@@
Figure 4.1. Supersonic regular reflection
Figure 4.2. Subsonic regular reflection
First, consider the problem with an assumption on the symmetry with respect to the $x_{1}$–axis. Then we can consider only the upper half-plane $\{x_{2}>0\}$ and prescribe the slip boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on the symmetry line $\{x_{2}=0\}$, so that there is only one reflection point $P_{0}$ to be considered (see Figs 4.24.2).
The regular shock reflection-diffraction configuration is characterized by the fact that the reflection occurs at point $P_{0}$ of the intersection of the incident shock with the wedge boundary. Figs. 4.24.2 show the structure of regular shock reflection-diffraction configurations in self-similar coordinates.
@@
Depending on the wedge angle, state $(2)$ can be either supersonic or subsonic at $P_{0}$, i.e., either $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ or the opposite inequality holds, where $c_{2}=\rho_{2}^{\frac{\gamma-1}{2}}$ is the (constant) speed of sound of state (2). Moreover, for $\theta_{\rm w}$ near $\frac{\pi}{2}$ (resp. for $\theta_{\rm w}$ near $\theta_{\rm w}^{\rm d}$), state $(2)$ is supersonic (resp. subsonic) at $P_{0}$. The type of state $(2)$ at $P_{0}$ for a given wedge angle $\theta_{\rm w}$ determines the type of reflection, supersonic or subsonic, as shown in Fig. 4.2 and Fig. 4.2, respectively.
@@
Next, we define the points and lines in Figs. 4.24.2. The incident shock $S_{0}$ is line $\{\xi_{1}=\xi_{1}^{0}\}$ with $\xi_{1}^{0}=\frac{\rho_{1}{u}^{(1)}_{1}}{\rho_{1}-\rho_{0}}>0$. The center, $O_{2}={\mathbf{u}}^{(2)}=(u_{1}^{(2)},u_{2}^{(2)})$, of the sonic circle $B_{c_{2}}(O_{2})$ of state $(2)$ lies on the wedge boundary between the reflection point $P_{0}$ and the wedge vertex $P_{3}$ for both the supersonic and subsonic cases.
Then, for the supersonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|>c_{2}$ so that $P_{0}\notin\overline{B_{c_{2}}(O_{2})}$, we denote by $P_{4}$ the upper point of intersection of $\partial B_{c_{2}}(O_{2})$ with the wedge boundary such that $O_{2}\in P_{3}P_{4}$. Also, the sonic circle $\partial B_{c_{2}}(O_{2})$ of state $(2)$ intersects line $S_{1}$, and one of the points of intersection, $P_{1}\in\Lambda$, is such that segment $P_{0}P_{1}$ is outside $B_{c_{2}}(O_{2})$. Denote the arc of $\partial B_{c_{2}}(O_{2})$ by $\Gamma_{\text{\rm sonic}}=P_{1}P_{4}$. The curved part of the reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{1}P_{2}$, where $P_{2}\in\{\xi_{2}=0\}$. Then we denote the line segments $\Gamma_{\rm sym}:=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}:=P_{3}P_{4}$. The lines and curves $\Gamma_{\text{\rm shock}}$, $\Gamma_{\text{\rm sonic}}$, $\Gamma_{\rm sym}$, and $\Gamma_{\text{\rm wedge}}$ do not have common points, except their endpoints $P_{1},\cdots,P_{4}$. Thus, $\Gamma_{\text{\rm shock}}\cup\Gamma_{\text{\rm sonic}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection. Denote by $\Omega$ the bounded domain restricted by this curve.
For the subsonic/sonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|\leq c_{2}$ so that $P_{0}\in\overline{B_{c_{2}}(O_{2})}$, the curved reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{0}P_{2}$, which does not have common interior points with the line segments $\Gamma_{\rm sym}=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}=P_{0}P_{3}$. Then $\Gamma_{\text{\rm shock}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection, and $\Omega$ is the bounded domain restricted by this curve.

### v2

@@ 2.2. Entropy Solutions and the Rankine-Hugoniot Conditions
$({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}+{\mathbf{v}}_{\varepsilon}+\nabla(h(\rho))_{\varepsilon}+\mathbf{r}^{(2)}_{\varepsilon}=0,$ (3.27)
@@
$r^{(1)}_{\varepsilon}=(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}),\qquad\mathbf{r}^{(2)}_{\varepsilon}=\big(({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big)_{\varepsilon}-({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}.$
@@
${\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}+(1+\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon})\omega_{\varepsilon}+\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}=0\qquad\,\,\mbox{in $\Omega^{r}$}.$ (3.28)
@@
$=\rho_{\varepsilon}X^{(\varepsilon)}+r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}$.
@@
${\mathcal{R}}_{\varepsilon}=f^{\prime}(X^{(\varepsilon)})\big(r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}\big)-f(X^{(\varepsilon)})r^{(1)}_{\varepsilon}$. (3.31)
@@
Let $\zeta\in C^{\infty}(\mathbb{R}^{2})$ satisfy (3.22). Multiply the last equation by $\zeta$, integrate over $\Omega$, and integrate by parts with the use of (3.22) and (3.25) to obtain
@@
[…]
@@
$\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}\to 0\qquad\,\mbox{in $L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.37)
@@
In the argument below, we use (3.24). We see that $\nabla(\rho_{\varepsilon}-\rho)\cdot{\mathbf{v}}_{\varepsilon}\to 0$ in $L^{2}(\Omega^{r})$ as $\varepsilon\to 0$, because $\nabla\rho_{\varepsilon}\to\nabla\rho$ in $L^{2}(\Omega^{r})$ and ${\mathbf{v}}_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$. Also, $(\rho_{\varepsilon}-\rho)(\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}-\mbox{\rm div}\,{\mathbf{v}})\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and $\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}\to\mbox{\rm div}\,{\mathbf{v}}$ in $L^{2}(\Omega^{r})$. Finally, $(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and converges to zero a.e. in $\Omega^{r}$, and $\mbox{\rm div}\,{\mathbf{v}}\in L^{2}(\Omega^{r})$, so that
@@
$\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}=\partial_{\xi_{1}}\big(({\mathbf{v}}\cdot\nabla v_{2})_{\varepsilon}-{\mathbf{v}}_{\varepsilon}\cdot\nabla(v_{2})_{\varepsilon}\big)-\partial_{\xi_{2}}\big(({\mathbf{v}}\cdot\nabla v_{1})_{\varepsilon}-{\mathbf{v}}_{\varepsilon}\cdot\nabla(v_{1})_{\varepsilon}\big).$
Then (3.37) follows from Lemma A.1 in Appendix A, applied in $\Omega^{\rm ext}_{r}$, with $p=q=2$, $b=v_{j}$, and $u=\partial_{j}v_{k}$ for the corresponding $j,k=1,2$, and i=3ki=3-k.
@@
where ${\mathbf{v}}^{-}=(v_{1}^{-},v_{2}^{-})=(u_{1}^{-}-\xi_{1},u_{2}^{-}-\xi_{2})$ with some constants $u_{1}^{-}$ and $u_{2}^{-}$, and ${\mathbf{v}}=(v_{1},v_{2})$. Taking the tangential derivative $\partial_{\boldsymbol{\tau}}:=\partial_{\xi_{1}}+f_{s}^{\prime}\partial_{\xi_{2}}$ of the Rankine-Hugoniot conditions (3.43) along $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and using the condition that $f_{s}^{\prime}=0$ at $P$, we obtain that, at point $P$,
@@
$0=-\int_{\Gamma_{\rm shock}}\big|\frac{\omega}{\rho}\big|^{2}\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}\,{\rm d}l>0.$
@@
This completes the proof of the fact that, under the assumptions of the theorem, it is not possible that (3.3) holds.
7. We now complete the proof as claimed. On the contrary, assume that
${\mathbf{v}}\in H^{1}(\Omega).$ (3.49)
Then, combining (3.49) with the assumptions in the theorem, we have
${\mathbf{v}}\in H^{1}(\Omega)\cap L^{\infty}(\Omega),\quad\,\,C^{-1}\leq\rho\leq C,\quad\,\,{\tilde{C}}^{-1}\leq p\leq\tilde{C}$, (3.50)
where $C,\;\tilde{C}\ \geq 1$, and we recall that $p=\frac{\rho^{\gamma}}{\gamma}$. We show that, with these properties, we can obtain from the weak form of the Euler system (2.6)–(2.7) that
$\int_{\Omega}\big((\rho{\mathbf{v}}\cdot\nabla{\mathbf{v}}+\rho{\mathbf{v}})\cdot\zeta-p\,\mbox{\rm div}\zeta\big)\,{\rm d}{\boldsymbol{\xi}}=0,$ (3.51)
for any $\zeta=(\zeta_{1},\zeta_{2})\in C^{1}_{\rm c}(\Omega;\mathbb{R}^{2})$. To show (3.51), we compute for $j=1,2$, using (3.50):
$\begin{split}\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}v_{j}\partial_{i}\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}&=\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}\partial_{i}(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}-\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}(\partial_{i}v_{j})\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}\\ &=\int_{\Omega}\rho{\mathbf{v}}\cdot\nabla(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}-\int_{\Omega}\rho({\mathbf{v}}\cdot\nabla v_{j})\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}.\end{split}$ (3.52)
Next, we note that, by the properties in (3.50), we can use the test function $\phi=v_{j}\zeta_{j}$ in (2.6) to obtain
$\int_{\Lambda}\rho{\mathbf{v}}\cdot\nabla(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}=\int_{\Omega}2\rho v_{j}\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}.$
Substituting this into the last expression of (3.52) for $j=1,2$, we have
$\int_{\Omega}\rho{\mathbf{v}}\otimes{\mathbf{v}}:D\zeta\,{\rm d}{\boldsymbol{\xi}}=\int_{\Omega}\big(2\rho v-\rho({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big)\cdot\zeta\,{\rm d}{\boldsymbol{\xi}}.$
Substituting this into (2.7), we obtain (3.51).
From (3.51) and the fact that $\rho{\mathbf{v}}\cdot\nabla{\mathbf{v}}+\rho{\mathbf{v}}\in L^{2}(\Omega)$ by (3.50), it follows from (3.51) that the weak derivative of $p$ exists in $\Omega$ and that $\nabla p\in L^{2}(\Omega)$. Then, using (3.50), we have
$p\in H^{1}(\Omega)\cap L^{\infty}(\Omega).$
From this, since $\rho=(\gamma p)^{1/\gamma}$ and ${\tilde{C}}^{-1}\leq p\leq\tilde{C}$, we obtain
$\rho\in H^{1}(\Omega)\cap L^{\infty}(\Omega).$
Combining this with (3.49), we obtain that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$, which contradicts (3.2). Then we conclude that (3.49) is false. ∎
@@ 4. Applications to Transonic Shock problems
Let $(\rho,{\mathbf{v}})$ be a solution of the Riemann problem that satisfies conditions (i)–(vii) of Definition 3.1 and assumption (i) of Theorem 3.2. Assume that
@@
    $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is not a segment of straight line. In particular, denoting by $P_{1}$ and $P_{2}$ the endpoints of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, there exists a point $P^{*}\in\overline{\Gamma_{1}^{\rm int}}\setminus\{P_{1}\}$ such that $\boldsymbol{\tau}(P_{1})\neq\pm\boldsymbol{\tau}(P^{*})$, where $\boldsymbol{\tau}(\cdot)$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at a point.
@@
Therefore, the remaining possibility is that $\hat{Q}=Q$. Note that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(Q)=0$. Moreover, using condition (c), the Rankine-Hugoniot conditions, and the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and $(\rho,{\mathbf{v}})$ given in condition (v) of Definition 3.1 and assumption (i) of Theorem 3.2, we see that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P_{1})=({\mathbf{v}}\cdot\boldsymbol{\tau}_{S})(P_{1})\neq 0$. This implies that $Q\neq P_{1}$. Also, since $\hat{Q}\neq P^{*}$ as we discussed above, then $Q\neq P^{*}$ for the present case. Thus, for the present case $Q=\hat{Q}$, it follows that $Q$ is an interior point of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, and the part of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ between $P_{1}$ and $Q$ lies on the straight line $S$. In particular, line $S$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q$. We now shift and rotate the coordinates to have the origin at $Q$ and the coordinates $\xi_{1}$ and $\xi_{2}$ to be along $S$ and $L$, respectively. Then $O_{1}=(0,\tilde{v}_{1})$ for some $\tilde{v}_{1}\in\mathbb{R}$ (in fact, $\tilde{v}_{1}\neq 0$ by condition (v) of Definition 3.1, but this will not be used below). To fix notation, let the $\xi_{1}$–axis along $S$ be oriented so that $P_{1}=(\xi_{P_{1}},0)$ with $\xi_{P_{1}}<0$. Since $S=\{\xi_{2}=0\}$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q=(0,0)$, then curve $\Gamma^{\mbox{\rm\footnotesize int}}_{1}\cap B_{r}(Q)$ is a graph for some $r>0$: There exists $f\in C^{2}(\mathbb{R})$ such that
@@
[…]
@@ 4.1. Lower Regularity of the Regular Shock Reflection Solutions for the Isentropic Euler System
$\Lambda=\mathbb{R}^{2}_{+}\setminus\big\{\mathbf{x}\;:\;x_{1}>0,\,0<x_{2}<x_{1}\tan\theta_{\rm w}\big\},$
where $\mathbb{R}^{2}_{+}=\mathbb{R}^{2}\cap\{x_{1}>0\}$. We seek functions $(\rho,{\mathbf{u}})(\mathbf{x},t)$ satisfying system (2.1) in $\Lambda$ with the boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on $\partial\Lambda$ and the initial data:
@@
Figure 2. Supersonic regular reflection
Figure 3. Subsonic regular reflection
First, consider the problem with an assumption on the symmetry with respect to the $x_{1}$–axis. Then we can consider only the upper half-plane $\{x_{2}>0\}$ and prescribe the slip boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on the symmetry line $\{x_{2}=0\}$, so that there is only one reflection point $P_{0}$ to be considered (see Figs 33).
The regular shock reflection-diffraction configuration is characterized by the fact that the reflection occurs at point $P_{0}$ of the intersection of the incident shock with the wedge boundary. Figs. 33 show the structure of regular shock reflection-diffraction configurations in self-similar coordinates.
@@
Depending on the wedge angle, state $(2)$ can be either supersonic or subsonic at $P_{0}$, i.e., either $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ or the opposite inequality holds, where $c_{2}=\rho_{2}^{\frac{\gamma-1}{2}}$ is the (constant) speed of sound of state (2). Moreover, for $\theta_{\rm w}$ near $\frac{\pi}{2}$ (resp. for $\theta_{\rm w}$ near $\theta_{\rm w}^{\rm d}$), state $(2)$ is supersonic (resp. subsonic) at $P_{0}$. The type of state $(2)$ at $P_{0}$ for a given wedge angle $\theta_{\rm w}$ determines the type of reflection, supersonic or subsonic, as shown in Fig. 3 and Fig. 3, respectively.
@@
Next, we define the points and lines in Figs. 33. The incident shock $S_{0}$ is line $\{\xi_{1}=\xi_{1}^{0}\}$ with $\xi_{1}^{0}=\frac{\rho_{1}{u}^{(1)}_{1}}{\rho_{1}-\rho_{0}}>0$. The center, $O_{2}={\mathbf{u}}^{(2)}=(u_{1}^{(2)},u_{2}^{(2)})$, of the sonic circle $B_{c_{2}}(O_{2})$ of state $(2)$ lies on the wedge boundary between the reflection point $P_{0}$ and the wedge vertex $P_{3}$ for both the supersonic and subsonic cases.
Then, for the supersonic case, i.e., when $|{\mathbf{v}}_{2}(P_{0})|=|P_{0}O_{2}|>c_{2}$ so that $P_{0}\notin\overline{B_{c_{2}}(O_{2})}$, we denote by $P_{4}$ the upper point of intersection of $\partial B_{c_{2}}(O_{2})$ with the wedge boundary such that $O_{2}\in P_{3}P_{4}$. Also, the sonic circle $\partial B_{c_{2}}(O_{2})$ of state $(2)$ intersects line $S_{1}$, and one of the points of intersection, $P_{1}\in\Lambda$, is such that segment $P_{0}P_{1}$ is outside $B_{c_{2}}(O_{2})$. Denote the arc of $\partial B_{c_{2}}(O_{2})$ by $\Gamma_{\text{\rm sonic}}=P_{1}P_{4}$. The curved part of the reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{1}P_{2}$, where $P_{2}\in\{\xi_{2}=0\}$. Then we denote the line segments $\Gamma_{\rm sym}:=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}:=P_{3}P_{4}$. The lines and curves $\Gamma_{\text{\rm shock}}$, $\Gamma_{\text{\rm sonic}}$, $\Gamma_{\rm sym}$, and $\Gamma_{\text{\rm wedge}}$ do not have common points, except for their endpoints $P_{1},\cdots,P_{4}$. Thus, $\Gamma_{\text{\rm shock}}\cup\Gamma_{\text{\rm sonic}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection. Denote by $\Omega$ the bounded domain restricted by this curve.
For the subsonic/sonic case, i.e., when $|{\mathbf{v}}_{2}(P_{0})|=|P_{0}O_{2}|\leq c_{2}$ so that $P_{0}\in\overline{B_{c_{2}}(O_{2})}$, the curved reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{0}P_{2}$, which does not have common interior points with the line segments $\Gamma_{\rm sym}=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}=P_{0}P_{3}$. Then $\Gamma_{\text{\rm shock}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection, and $\Omega$ is the bounded domain restricted by this curve.